OpenStudy (anonymous):

Which is a rational number? A. 3.1 repeating B. pi C. √2 D. √3

OpenStudy (igreen):

Rational numbers are numbers that can be written as simple fractions.

OpenStudy (igreen):

So what do you think the answer is?

OpenStudy (anonymous):

i think its c

OpenStudy (anonymous):

So it has to be C or D right?

OpenStudy (igreen):

@XsanayX Don't give answers, it's against the code of conduct.

OpenStudy (igreen):

Actually..it's none of those.

OpenStudy (michele_laino):

I think 3.1 since I can write it as a fraction, namely: \[3.1=\frac{ 31 }{ 10 }\] other numbers can not be expressed by a fraction, in other words can not exist two integer numbers, say m and n such that: \[\frac{ m }{ n }=\pi\] or \[\frac{ m }{ n }=\sqrt{2}\] or: \[\frac{ m }{ n }=\sqrt{3}\]

OpenStudy (igreen):

@Michele_Laino You're not supposed to give answers..we're here to guide the asker to a solution so that the asker may get it on their own.

OpenStudy (igreen):

Plus it's not 3.1, it's \(3. \overline{1}\)

OpenStudy (igreen):

Which means..you don't give the answer at all..no matter if you give an explanation or not.

OpenStudy (michele_laino):

ok! you are right! @iGreen Please @Vivid my answer is incorrect. Your correct answer is the one that @iGreen gave you before, namely none of your numbers that you provided is a rational number

OpenStudy (michele_laino):

please, the right explanation is Always a benefit @iGreen

OpenStudy (michele_laino):

oops...is always...

OpenStudy (michele_laino):

I'm very sorry for my incorrect answer @Vivid

OpenStudy (igreen):

Well..your answer is wrong because '3.1' isn't even a choice..but however \(3. \overline{1}\) was the correct answer..and he already walked off with that choice.

OpenStudy (igreen):

Yes..but the problem is the asker could just take your answer and not read your explanation at all..

OpenStudy (michele_laino):

yes! the correct answer is C thanks for your advice @iGreen